# GHP-MOFassemble: Diffusion modeling, high throughput screening, and molecular dynamics for rational discovery of novel metal-organic frameworks for carbon capture at scale

### v1

@@ Abstract
We introduce GHP-MOFassemble, a Generative artificial intelligence (AI), High Performance framework to accelerate the rational design of metal-organic frameworks (MOFs) with high CO2 capacity and synthesizable linkers. Our framework combines a diffusion model, a class of generative AI, to generate novel linkers that are assembled with one of three pre-selected nodes into MOFs in a primitive cubic (pcu) topology. The CO2 capacities of these AI-generated MOFs are predicted using a modified version of the crystal graph convolutional neural network model. We then use the LAMMPS code to perform molecular dynamics simulations to relax the AI-generated MOF structures, and identify those that converge to stable structures, and maintain their porous properties throughout the simulations. Among 120,000 pcu MOF candidates generated by the GHP-MOFassemble framework, with three distinct metal nodes (Cu paddlewheel, Zn paddlewheel, Zn tetramer), a total of 102 structures completed molecular dynamics simulations at 1 bar with predicted CO2 capacity higher than 2 mmol/g at 0.1 bar, which corresponds to the top 5% of hMOFs in the hypothetical MOF (hMOF) dataset in the MOFX-DB database. Among these candidates, 18 have change in density lower than 1% during molecular dynamics simulations, indicating their stability. We also found that the top five GHP-MOFassemble’s MOF structures have CO2 capacities higher than 96.9% of hMOF structures. This new approach combines generative AI, graph modeling, large-scale molecular dynamics simulations, and extreme scale computing to open up new pathways for the accelerated discovery of novel MOF structures at scale.
  • July 13, 2023
Keywords: diffusion model, pre-trained regression model, metal-organic framework, carbon capture
1 Introduction
@@
MOFs have been shown to exhibit superior chemical and physical CO2 adsorption properties. They can be recycled in appropriate operating environments for varying numbers of times before undergoing significant structural degradation. However, their industrial applications have not yet reached full potential due to stability issues, such as poor long-term recyclability [5], and high moisture sensitivity [6]. For instance, the presence of moisture in adsorption gas impairs a MOF’s CO2 capture performance, which may be attributed to MOFs having stronger affinity toward water molecules than to CO2 molecules [7]. Other investigations of MOF stability issues [8, 9, 10] have shown [11, 12] that the gas adsorption properties of MOFs can be enhanced by tuning the building blocks used. Yet progress has been difficult due to the enormous chemical space of building blocks that makes exhaustive search with traditional experimental methods impractical [13].
Here we propose GHP-MOFassemble, a novel high-throughput computational framework to accelerate the discovery of MOF structures with high CO2 capacities and synthesizable linkers. To demonstrate the utility of our framework, we apply it to the rational design of MOFs with pcu topology and three types of inorganic nodes: Cu paddlewheel, Zn paddlewheel, and Zn tetramer. The key component of our generative framework is a diffusion model, Difflinker [14], which enables efficient generation of new linker molecules from a set of molecular fragments.
Previous attempts to discover useful MOFs have proceeded via high-throughput screening of existing datasets [15, 16, 17], an approach that necessarily limits the search to known MOFs. Our GHP-MOFassemble framework instead probes the MOF design space by employing a molecular generative diffusion model (DiffLinker [14]) to generate chemically diverse and unique MOF linkers, which it assembles with pre-selected metal nodes to form novel MOF structures. It then screens those structures with a pre-trained regression model, a modified version of CGCNN [18], to identify high-performing MOF candidates for carbon capture.
2 Related Work
@@
Previous efforts in the search for new MOF structures with exceptional gas adsorption properties include high-throughput screening and generative modeling. High-throughput screening approaches comprise both database search and machine learning (ML)–assisted screening.
@@
Database search methods. These methods apply filters to identify attractive candidates in large databases of MOF structures plus calculated properties obtained by using molecular simulations. For example, Li et al. [16] search the computation-ready, experimental MOF database (CoRE DB) for MOFs with high $\textrm{CO}_{2}$ uptake when moisture is present, and Altintas et al. [19] search both CoRE DB and the Cambridge Structural Database non-disordered MOF subset for MOFs with high $\textrm{CH}_{4}/\textrm{H}_{2}$ selectivity.
@@
ML–assisted screening. Database search methods require expensive molecular simulation calculations for every target molecule. ML-assisted screening avoids this difficulty by using a regression model trained on a smaller training dataset to predict target properties for a large number of new test structures. This approach has been widely applied to gas adsorption and separation property predictions of MOFs, including CO2/H2 separation [20] and CH4 adsorption [21]. ML-assisted methods commonly make use of both geometrical features (e.g., dominant pore size, void fraction [22]) and chemical features (e.g., atom type, electronegativity [23]). A compound feature, the atomic property weighted radial distribution function (AP-RDF) [24] has been shown to improve regression model performance in finding MOF structures with high CO2 capacity [25]. This feature is constructed by using both chemistry and local geometry of atomic sites in MOF structures. As an alternative to the use of hand-engineered machine learning features, neural networks have also been applied to MOF research. In a recent study, the Atomistic Line Graph Neural Network (ALIGNN), trained on the hMOF dataset [26], was applied to search CoRE DB for high-performing MOF candidates for carbon capture [27].
@@
Generative modeling. In the generative modeling approach, candidate molecules are not drawn from a database but are generated de novo via methods that produce novel compounds that are expected to have a desired set of chemical features. One notable example is the recently proposed Supramolecular Variational Autoencoder (SmVAE[28], which uses a semantically constrained graph-based canonical code (RFcode) to encode MOF building blocks. The variational autoencoder framework structure allows this model to interpolate between existing MOF structures and enables isoreticular optimization of MOF structures toward higher CO2 capacity and CO2/N2 selectivity.
@@
This work. Proposed generative models include not only variational autoencoders (VAEs) but also generative adversarial networks (GANs), normalizing flows, autoregressive models, and diffusion models [29]. In this work, we adopt a diffusion model named DiffLinker to generate novel MOF linkers. Diffusion models use a probability distribution and Markovian properties to generate new data via a denoising step. First, Gaussian noise is added to the input samples to yield noisy data. A neural network is then trained to learn what noise was added. The trained network is then used to reversely transform (i.e., denoise) the noisy data back into target samples that resemble molecules from the training data distribution. This method has been widely applied to drug discovery to speed up the design of new ligands and ligand-protein docking configurations [30, 31, 32, 33].
Two broad categories of molecular representation schemes have been used in diffusion model-based ligand generators: molecular graphs and 3D coordinates. In the former case (Digress/Congress [34] is an example), atoms are represented as nodes and bonds as edges. In the latter case, the 3D atomic coordinates of molecules are generated directly, as in DiffLinker [14] and E(3) equivariant diffusion model (EDM[35]. Given the success of diffusion models in drug discovery [32, 30, 33, 36, 37], in Section 3.1 we demonstrate how to transfer the idea to the iso-reticular design of MOF structures by varying MOF linkers while fixing node and topology. The diffusion model is used specifically for MOF linker design, as we describe in Section 3.2.
3 Methods
@@
  • Decompose. We use a molecular fragmentation algorithm to decompose the MOF linkers found in high-performing MOFs within the hMOF dataset [26]—an open source dataset that provides, for each of 137,652 hypothetical MOF structures, its MOFid, MOFkey, geometric features, and isotherm data for six adsorption gases (CO2, N2, CH4, H2, Kr, Xe) at 0.01, 0.05, 0.1, 0.5, and 2.5 bar—covering the pressure ranges of cyclic adsorption gas separation processes in industrial applications. The adsorption properties of gases provided in this dataset were calculated using Grand Canonical Monte Carlo (GCMC) calculations [26], as described in Wilmer et al. [38].
@@
  • Generate. We use a diffusion model to generate new MOF linkers. We then screen the generated linkers by removing linkers with S, Br and I elements (we call this step “element filter” in later text), and evaluate their quality using five different scores to quantify their synthesizability, validity, uniqueness, and internal diversity. Linkers that passed the element filter are then assembled with one of three pre-selected nodes into new MOF structures in the pcu topology.
  • Predict. We use a pre-trained regression model to predict the CO2 capacities of the newly generated MOF structures.
@@
In the following sections we describe each of these components in detail.
3.1 Decomposing MOF linkers into molecular fragments
@@
The first component of the GHP-MOFassemble framework, Decompose, decomposes linkers from high-performing MOF structures in the hMOF dataset into their molecular fragments. As illustrated in Figure 1, this process applies the following three steps to a specified node-topology pair, which in the figure is the most frequent node-topology pair in the hMOF dataset, Zn tetramer-pcu. Note that for the pcu topology, one linker is orientated along each of the x, y, and z directions, and thus at most three linker types are possible.
@@
    Select: We select high-performing MOFs with the given node-topology pair from hMOF. We define here a high-performing MOF structure as one with CO2 capacity higher than 2 mmol/g at 0.1 bar, which corresponds to the top 5% of CO2 capacity.
@@
    Extract: We extract the linkers Simplified Molecular-Input Line-Entry System (SMILES) strings for each high-performing MOF identified in step (i), eliminate those for MOFs with more than three unique linker types, and assemble those that remain into a linker dataset.

### v2

# A generative artificial intelligence framework based on a molecular diffusion model for the design of metal–organic frameworks for carbon capture

@@ Abstract
Metal–organic frameworks (MOFs) exhibit great promise for CO2 capture. However, finding the best performing materials poses computational and experimental grand challenges in view of the vast chemical space of potential building blocks. Here, we introduce GHP-MOFassemble, a generative artificial intelligence (AI), high performance framework for the rational and accelerated design of MOFs with high CO2 adsorption capacity and synthesizable linkers. GHP-MOFassemble generates novel linkers, assembled with one of three pre-selected metal nodes (Cu paddlewheel, Zn paddlewheel, Zn tetramer) into MOFs in a primitive cubic topology. GHP-MOFassemble screens and validates AI-generated MOFs for uniqueness, synthesizability, structural validity, uses molecular dynamics simulations to study their stability and chemical consistency, and crystal graph neural networks and Grand Canonical Monte Carlo simulations to quantify their CO2 adsorption capacities. We present the top six AI-generated MOFs with CO2 capacities greater than 2m $\textrm{m\,mol}\,\textrm{g}^{-1}$, i.e., higher than 96.9% of structures in the hypothetical MOF dataset.
Introduction
@@
MOFs have been shown to exhibit superior chemical and physical CO2 adsorption properties. In appropriate operating environments, they can be recycled, for varying numbers of times, before undergoing significant structural degradation. However, their industrial applications have not yet reached full potential due to stability issues, such as poor long-term recyclability [5], and high moisture sensitivity [6]. For instance, the presence of moisture in adsorption gas impairs a MOF’s CO2 capture performance, which may be attributed to MOFs having stronger affinity toward water molecules than to CO2 molecules [7]. Other investigations of MOF stability issues [8, 9, 10, 11, 12] have shown that the gas adsorption properties of MOFs can be enhanced by tuning the building blocks used. Yet progress has been difficult due to the enormous chemical space of building blocks, which makes exhaustive search with traditional experimental methods impractical [13].
Related Work
@@
Previous efforts in the search for new MOF structures with exceptional gas adsorption properties include:
@@
Database search methods. These methods apply filters to identify optimal candidates in large databases of MOF structures plus calculated properties obtained with molecular simulations. For example, Ref. [14] shows how to search for MOFs with high $\textrm{CO}_{2}$ uptake when moisture is present using the experimental MOF database (CoRE DB), while Ref. [15] shows how to search for MOFs with high $\textrm{CH}_{4}/\textrm{H}_{2}$ selectivity in both CoRE DB and the Cambridge Structural Database non-disordered MOF subset.
@@
Machine Learning (ML)-assisted screening. Building large MOF databases require expensive molecular simulation calculations for every target structure. ML-assisted screening avoids this difficulty by using a regression model trained on a smaller training set to predict target properties of a large number of new test structures. This approach has been widely applied to gas adsorption and separation property predictions of MOFs, including CO2/H2 separation [16] and CH4 adsorption [17]. ML-assisted methods use of geometrical features (e.g., dominant pore size, void fraction [18]) and chemical features (e.g., atom type, electronegativity [19]). A compound feature, the atomic property weighted radial distribution function [20] has been shown to improve regression model performance in finding MOF structures with high CO2 capacity [21]. This feature is constructed by using both local geometry and chemistry of atomic sites in MOF structures. As an alternative to the use of hand-engineered ML features, neural networks have been applied to MOF research. In Ref. [22], the Atomistic Line Graph Neural Network, trained on the hMOF dataset, was applied to search CoRE DB for high-performing MOFs for carbon capture [23].
@@
Generative modeling. Candidates are not drawn from a database but are generated de novo via methods that produce novel compounds that have a desired set of chemical features. For instance, the Supramolecular Variational Autoencoder [24] uses a semantically constrained graph-based canonical code to encode MOF building blocks. The variational autoencoder framework structure allows this model to interpolate between existing MOF structures and enables isoreticular optimization of MOF structures toward higher CO2 capacity and CO2/N2 selectivity.
@@
This work. Proposed generative models include variational autoencoders, generative adversarial networks, normalizing flows, autoregressive models, and diffusion models [25]. Here, we adopt a diffusion model named DiffLinker to generate novel MOF linkers. Diffusion models use a probability distribution and Markovian properties to generate new data via forward diffusion and backward denoising steps. First, Gaussian noise is added to the input samples to yield noisy data. A neural network is then trained to learn what noise was added. The trained network is then used to reversely transform (i.e., denoise) the noisy data back into target samples that resemble molecules from the training data distribution. This method has been widely applied to drug discovery to speed up the design of new ligands and ligand-protein complexes [26, 27, 28, 29]. Two broad categories of molecular representation schemes have been used in diffusion model-based ligand generators: molecular graphs and 3D coordinates. In the former case (Digress/Congress [30] is an example), atoms are represented as nodes and bonds as edges. In the latter case, the 3D atomic coordinates of molecules are generated directly, as in DiffLinker [31] and E(3) equivariant diffusion model (EDM) [32]. Given the success of diffusion models in drug discovery [28, 26, 29, 33, 34], we demonstrate how to transfer the idea to the iso-reticular design of MOF structures by varying MOF linkers while fixing node and topology. The diffusion model is used specifically for MOF linker design.
Our proposed approach, GHP-MOFassemble, is a novel high-throughput computational framework to accelerate the discovery of MOF structures with high CO2 capacities and synthesizable linkers. While previous attempts to discover useful MOFs via computational methods have proceeded via high-throughput screening of existing datasets [35, 14, 36], an approach that necessarily limits the search to known MOFs. In contrast, GHP-MOFassemble probes the MOF design space by employing a molecular generative diffusion model, DiffLinker [31], to generate chemically diverse and unique MOF linkers from a set of molecular fragments, which it assembles with pre-selected metal nodes to form novel MOF structures. It then screens those structures with a pre-trained regression model, a modified version of Crystal Graph Convolutional Neural Network (CGCNN) [37], to identify high-performing MOF candidates for carbon capture. In addition, we apply molecular dynamics (MD) and grand canonical Monte Carlo (GCMC) to further down-select stable AI-generated MOFs and compute more credible CO2 adsorption capabilities. We demonstrate the utility of our framework by applying it to the rational design of MOFs with pcu topology and three types of inorganic nodes: Cu paddlewheel, Zn paddlewheel, and Zn tetramer.
Results
Here we describe the key components of GHP-MOFassemble, and present a new set of AI-generated MOF structures with high CO2 capacities. A detailed description of the approaches used to obtain these findings is provided in Methods.
Analysis of the hMOF dataset
The three most frequent node-topology pairs in the hMOF dataset, accounting for around 74% of its MOFs, are the Cu paddlewheel-pcu, Zn paddlewheel-pcu, and Zn tetramer-pcu, which amount to 102,117 hMOF structures (29,714 + 28,529 + 43,874 from first column of Table 1). Out of these 102,117 structures, we only used those with correctly parsed MOFids and valid Simplified Molecular-Input Line-Entry System (SMILES), i.e., 78,238 MOFs.
Table 1: Most frequent node-topology pairs in the hMOF dataset. Properties of the hMOF dataset. PW, TM, and HP-MOF stand for paddlewheel, tetramer, and high-performing MOF, respectively.
three linkers SMILES fragment conformers
Cu PW-pcu 29,714 1,458 1,016 3,330 180
Zn PW-pcu 28,529 1,314 834 3,221 162
Zn TM-pcu 43,874 2,129 1,388 3,265 198
Linker generation and evaluation
For all three MOF types, we start with 540 (i.e., = 180 + 162 + 198 from Table 1 last column) unique molecular fragments extracted from high-performing hMOF structures, and use DiffLinker to generate new MOF linkers, with the number of sampled connection atoms ranging from five to 10 (total six connection atom types). For each linker, sampling is performed 20 times. Therefore, the resulting candidate pool contains a total of 64,800 linkers, i.e., 540 fragments $\times$ 20 independent sampling steps $\times$ six unique number of connection atoms. Since these linkers only contain heavy atoms, to generate all-atom linker molecules, we apply openbabel to add hydrogen atoms, which results in 56,257 linkers after removing linkers with erroneous hydrogen assignments. Next, dummy atom identification is performed to generate information that enables assembly with metal nodes. A total of 16,162 linkers with dummy atoms are generated. These linkers are then passed through the element filter, which removes linkers that contain S, Br, and I, further reducing the number to 12,305 linkers. The number of linkers in each step are summarized in the Total column of Table 2.
Table 2: Number of linkers at each step of the MOF assembly process. Statistical summary of the number of linkers after each step categorized by the corresponding MOF types. PW and TM stand for paddlewheel and tetramer, respectively. Element filter means that linkers with S, Br and I elements are removed.
Method Total Cu PW-pcu Zn PW-pcu Zn TM-pcu
DiffLinker 64,800 21,600 19,440 23,760
Hydrogen addition 56,257 18,979 17,126 20,152
Dummy atom identification 16,162 4,964 4,450 6,748
Element filter 12,305 3,702 3,441 5,162
MOF assembly
We generate new MOF structures by assembling three randomly selected DiffLinker-generated linkers with one of the three most frequent nodes in the hMOF dataset. Random sampling of 12,305 linkers ensures that the selected linkers cover a large design space. For each node-linker-topology combination, we considered four levels of catenation (cat0, cat1, cat2, cat3). We generated a total of 120,000 MOFs as follows: we have four catenation levels and three node candidates. The random sampling of 10,000 linkers for each catenation level-node candidate pair generates 120,000 total MOFs of different catenation-node-linker combinations.
MOF geometry examination
Inter-atomic distance check
We used Pymatgen to read each MOF structure’s CIF file and to extract the pairwise distance matrix. Diagonal entries of the distance matrix are discarded as they signify the distance between an atom and itself. The off-diagonal values are tested for minimum inter-atomic distance threshold. The inter-atomic distance threshold is predetermined by an experimental database, OChemDb [38]. If a MOF has at least one entry of lower than threshold inter-atomic distance, the MOF is discarded. 78,796 of the 120,000 assembled MOFs passed this test.
Pre-simulation Check
We used the open source library, cif2lammps [39], to automatically assign the Universal Force Field for Metal–Organic Frameworks (UFF4MOF) [40, 41] to MOFs and generate LAMMPS [42] input files. This step ensures that all atomic structures and bonding appearing in each MOF structure are chemically valid within the scope of UFF4MOF. For example, if a Zn atom is bonded to a C atom, or if an O atom is bonded to four other atoms, etc., these structures will be identified as invalid and will be discarded. Of the 78,796 MOFs that passed geometry and inter-atomic distance checks, 18,770 passed this pre-simulation check, and thus had LAMMPS input files generated.
Regression model for MOF $\textrm{CO}_{2}$ capacity prediction
To reduce the number of LAMMPS simulations, a screening step of the MOFs’ adsorption performance is conducted on the 18,770 MOFs described above using a modified version of the CGCNN model that we introduced in Ref. [37]. Here, we used MOF structures from the hMOF dataset as well as their CO2 capacities at 0.1 bar as input data. We split the hMOF dataset into three independent sets: 80% for training, 10% for validation, and 10% for testing. Using this data split, we trained three CGCNN models using random initialization of weights for each of them. When we use this model ensemble to infer the CO2 capacities of newly AI-generated MOF structures, we take the average of the predictions made by the three independent models as the predicted CO2 adsorption capacity. Out of the 18,770 AI-generated MOF structures, a total of 364 were predicted to be high-performing MOFs with CO2 capacity higher than 2 m $\textrm{m\,mol}\,\textrm{g}^{-1}$ at 0.1 bar. Specifically, we used ensemble model prediction mean plus standard deviation as CO2 adsorption capacity value to be higher than 2 m $\textrm{m\,mol}\,\textrm{g}^{-1}$ as the threshold. The standard deviation consideration is to assure we account for statistical errors of adsorption predictions by our ensemble model. Categorization of the predicted high-performing MOFs by node-topology pair and catenation level indicates that for all three node types most of the high-performing MOFs are cat2 and cat3.
Structural validation of AI-generated MOFs with molecular dynamics simulations
We now examine the stability and porous properties of the 364 AI-generated MOFs described in the previous section using MD simulations with LAMMPS [42]. For each MOF, a 2×2×2 supercell structure is equilibrated under a triclinic isothermal-isobaric ensemble (i.e., NPT) at $\langle p\rangle=1\ \textrm{atm}$ and $\langle T\rangle=300\textrm{K}$, such that the cell lengths and angles of the MOF structure can be equilibrated. The NPT simulations are run for 400,000 steps with step size of 0.5 femtoseconds. The relative changes of the lattice parameters before and after the simulation are inspected to evaluate how well the MOF structure is maintained throughout the equilibrium MD simulation. The MOFs with higher than 5% changes in any of the lattice parameters ($a$, $b$, $c$, $\alpha$, $\beta$, $\gamma$) are discarded (first three are lattice vector lengths and the latter three are lattice angles) between the initial state and the equilibrated state. The MOFs with higher than 5% changes in any of the lattice parameters ($a$, $b$, $c$, $\alpha$, $\beta$, $\gamma$) are discarded. This process produces 102 MOFs that satisfied the previously defined criteria, i.e., $<5\%$ in lattice parameter change [40].
Property validation of AI-generated MOFs with grand canonical Monte Carlo (GCMC) simulations
These 102 MOFs are further examined with GCMC simulations to calculate their CO2 adsorption capacity at a pressure and temperature of 0.1 bar and 300 K, respectively. RASPA [43] is used to conduct these GCMC CO2 adsorption simulations. As demonstrated in Figure 1, 6 MOFs candidates were found to have CO2 adsorption capacity higher than 2 m $\textrm{m\,mol}\,\textrm{g}^{-1}$ by the GCMC simulations. The 3D structure of these six high performing, AI-generated MOFs is presented in Figure 2.
Figure 1: CO2 adsorption values at 0.1 bar and 300K. CO2 adsorption capacity values at a pressure and temperature of 0.1 bar and 300 K, respectively, for the top six AI-generated MOFs’ according to grand canonical Monte Carlo (GCMC) simulations and our modified crystal graph convolutional neural network (CGCNN) model.
Figure 2: Visualization of the crystal structure of AI-generated MOFs. a–f Crystal structure of the top six AI-generated MOFs. The color code used to represent atoms is: carbon in grey, nitrogen in dark blue, fluorine in cyan, zinc in purple, hydrogen in white, and lithium in green.
Discussion
GHP-MOFassemble combines novel generative and graph AI applications, as well as a comprehensive screening workflow that combines various modeling methods with increasing levels of chemistry awareness and increasing levels of computational cost. When we deployed and optimized GHP-MOFassemble on the Delta supercomputer, unless indicated otherwise, we found that:
  • We AI-assembled 120,000 MOFs within 33 minutes using multiprocessing on 28 cores in the ThetaKNL supercomputer at the Argonne Leadership Computing Facility (ALCF).
  • We screened these 120,000 MOFs in 40 minutes using multiprocessing on 128 cores, and identified 78,796 MOFs with valid bond lengths.
  • We further screened these 78,796 MOFs and identified 18,770 MOFs with valid chemistry, using UFF4MOF as a reference, within 205 minutes using multiprocessing on 128 cores.
  • We then used our AI ensemble of CGCNN models to estimate the CO2 capacity of these 18,770 MOFs, and identified 364 high-performing, AI-generated MOFs with CO2 capacity higher than 2 m $\textrm{m\,mol}\,\textrm{g}^{-1}$ at 0.1 bar. AI inference was completed in 50 minutes using one NVIDIA A40 GPU.
In brief, from assembly to selection of high-performing MOFs, GHP-MOFassemble completes the analysis within 5 hours and 7 minutes. Once we have selected 364 AI-generated, high-performing MOFs, we carry out the most compute-intensive part of the analysis, namely:
  • We used the LAMMPS code to equilibrate 364 MOFs with 200-picosecond NPT MD simulations with the UFF4MOF force field. Each of these simulations is completed within 11 minutes, on average, using between 6 to 14 MPI processes, depending on the number of atoms in the MOF. We identified 102 stable MOFs whose lattice parameters changed less than 5% throughout the MD simulations.
  • Finally, each of these 102 MOFs were further examined with GCMC simulations, and a total of six MOF candidates were found to have CO2 capacity higher than 2 m $\textrm{m\,mol}\,\textrm{g}^{-1}$, which corresponds to the top 5% of the hMOF dataset. Each of these 102 GCMC simulations is completed within six hours, on average, using one CPU core.
Recent studies [44, 45, 46, 47] indicate that several functional groups play an important role in determining MOF CO2 capacity, including carboxylic acid (-COOH), primary amine (-NH2), hydroxyl (-OH), and nitrile (-CN). These functional groups affect MOF CO2 capacity due to their interaction with CO2 molecules through charge redistribution [46], hydrogen bond interaction [45], and electrostatic interaction [45]. Thus, we analyzed the functional groups of linkers in the 364 AI-generated high-performing MOF structures and compared them with high-performing hMOF structures. The results of this analysis are presented in Figure 3. We observe that linkers in high-performing hMOF structures have a higher proportion of carboxylic groups, whereas hydroxyl groups appear more often in the predicted high-performing MOFs generated by our framework. Moreover, the percentages of primary amine and nitrile in linkers are similar in both cases. Note that some substructures appear more often than others. For example, ring structures appear in most of the linkers in the top candidates, and many of the DiffLinker-generated molecular substructures (in between the terminal fragments) also contain ring structures. The frequent occurrence of rings in linkers of high-performing MOFs may be due to the presence of CO2-ring interaction, as reported in Ref. [48]. In addition, rings in the linkers may increase MOFs’ structural rigidity due to less rotatable bonds.
In summary, the entire discovery process from generative AI MOF assembly to detailed GCMC simulations of top AI-predicted MOFs can be completed within 12 hours using distributed computing in modern supercomputing platforms. The entire workflow may be further accelerated by scaling up any of the subprocesses using more cores or by distributing AI inference over more GPUs. Furthermore, the various pieces of GHP-MOFassemble may be assembled to create a standalone workflow that combines MD, density functional theory, and GCMC simulations to expose our generative AI model to a larger set of chemically diverse, high performing MOFs. Through online learning methods one may continually guide generative AI until it consistently assembles high performing MOFs. A study of this nature will be pursued in future work.
Figure 3: Functional groups in high-performing MOFs. Comparison of the proportion of selected functional group that appear in high-performing AI-generated MOFs and MOFs in the hMOF dataset.
Methods
@@
  • Decompose. We use a molecular fragmentation algorithm to decompose the MOF linkers found in high-performing MOF structures within the hMOF dataset [22]—an open source dataset that provides, for each of 137,652 hypothetical MOF structures, and corresponding MOFid, MOFkey, geometric features, and isotherm data for six adsorption gases (CO2, N2, CH4, H2, Kr, Xe) at 0.01, 0.05, 0.1, 0.5, and 2.5 bar—covering the pressure ranges of cyclic adsorption gas separation processes in industrial applications. The adsorption properties of gases provided in this dataset were calculated using GCMC calculations [22], as described in Ref. [49].
@@
  • Generate. We use a diffusion model to generate new MOF linkers. We then screen the AI-generated linkers by removing linkers with S, Br and I elements (we call this step “element filter”), and evaluate their quality using five different scores to quantify their synthesizability (SAScore and SCScore), validity, uniqueness, and internal diversity. Linkers that passed the element filter are then assembled with one of three pre-selected nodes into new MOF structures in the pcu topology.
  • Screen and Predict. We check for inter-atomic distances with pre-simulation checks to ensure we filter poor AI-generated MOF structures from previous steps. We then use a pre-trained regression model to predict the CO2 capacities of the newly generated MOF structures. Lastly, we perform MD and GCMC simulations to obtain stable and high-performing MOF structures.
@@
In the following sections we describe the datasets we have used, and each of the GHP-MOFassemble components in detail.
The hMOF dataset
We selected the most common topologies of the hMOF dataset for this work, i.e., Cu paddlewheel-pcu, Zn paddlewheel-pcu, and Zn tetramer-pcu, i.e., 78,238 MOFs with correctly parsed MOFids and valid SMILES. We gather the numbers of molecular fragments in Table 1, produced through the Fragment step of our GHP-MOFassemble framework, described below. In Table 1 high-performing MOFs (second column) with three parsed linkers are selected (third column), and their unique linkers are parsed by using the MMPA (Matched Molecular Pairs Algorithm) algorithm. The cumulative distribution functions of the CO2 capacities of these three types of MOFs are shown in the right panel of Figure 4.
Figure 4: Properties of the hMOF dataset. a Depictions of the most frequent node-topology pairs in hMOF structures. b Cumulative distribution functions of their 0.1 bar CO2 capacities.
In Table 1, the number of output molecular fragment conformers (last column) is much less than the number of unique linker SMILES (second to last column) for two reasons. First, around 56% of linkers did not pass the valency check, which may be due to the intrinsic limitations in the parsing of SMILES of MOF linkers. Second, around 90% of linkers that passed the valency check share similar molecular fragments, and thus many duplicated molecular fragments exist. The successfully parsed unique molecular fragment conformers are subsequently used for linker generation.
Figure 5 shows the empirical cumulative distribution functions of the CO2 capacities of hMOF structures at different catenation levels (i.e., MOFs with interpenetrated lattices). Therein, we observe that a higher percentage of catenated MOFs (cat1, cat2 and cat3) are high-performing, as compared to the uncatenated MOFs (cat0). This result confirms that catenation is an important factor when designing new MOF structures with high CO2 capacity. This observation is consistent with other studies in the literature [50, 51], which indicate that even though catenation reduces pore size and surface areas, catenated MOFs generally have higher CO2/H2 selectivities because MOF-CO2 interactions are enhanced as a result of the strong confinement of CO2 with a much lower adsorption surface. Thus, the results presented in Figure 5 using the hMOF dataset indicate that CO2 working capacities of catenated MOFs are higher than their noncatenated counterparts. As we discuss below, we found a similar pattern in AI-generated MOFs. Catenated MOFs were generated by using site translation method, which is achieved by displacing the reference lattice along the diagonal line of the unit cell. For all four catenation levels, the amount of relative lattice displacement is given in Table 3. The numbers are fractional displacements relative to the unit cell diagonal.
Figure 5: CO2 capacities of hMOF structures at different catenation levels. Empirical cumulative distribution functions of MOFs in the hMOF dataset at 0.1 bar at different catenation levels. The x axis is capped at 4 m $\textrm{m\,mol}\,\textrm{g}^{-1}$ to preserve details.
Table 3: Relative lattice displacement at each catenation level. Amount of lattice displacement for catenated MOFs.
catenation lattice1 lattice2 lattice3 lattice4
cat0 (0,0,0) / / /
cat1 (0,0,0) ($\frac{1}{2}$,$\frac{1}{2}$,$\frac{1}{2}$) / /
Figure 6 shows the pairwise relationships of the CO2 capacities of hMOF structures at five pressures. We observe a strong correlation of CO2 capacities between 0.05 bar and 0.1 bar, with a Pearson’s correlation coefficient of 0.86. For other pairs of pressures, the CO2 capacities are only weakly correlated, with a decreasing correlation for larger pressure differences. Moreover, the distributions of CO2 capacities at all pressures exhibit long tails at high value ranges, which indicate that the majority of MOFs are low-performing and high-performing MOFs are uncommon.
Figure 6: $\textrm{CO}_{2}$ capacities of hMOF structures at different adsorption pressures. Pairplot of $\textrm{CO}_{2}$ capacities of hMOF structures. The $(x,y)$ axes represent adsorption pressures. Different colors indicate MOFs with different node-topology pairs.
Decomposing MOF linkers into molecular fragments
@@
The first component of the GHP-MOFassemble framework, Decompose, decomposes linkers from high-performing MOF structures in the hMOF dataset into their molecular fragments. As illustrated in Figure 7, this process applies the following three steps to a specified node-topology pair, which in the figure is the most frequent node-topology pair in the hMOF dataset, Zn tetramer-pcu. Note that for the pcu topology, one linker is orientated along each of the x, y, and z directions, and thus at most three unique linker types are possible.
@@
    Select: We select high-performing MOFs with the given node-topology pair from hMOF. We define here a high-performing MOF structure as one with CO2 capacity higher than 2 m $\textrm{m\,mol}\,\textrm{g}^{-1}$ at 0.1 bar, which corresponds to the top 5% of CO2 capacity of the hMOF dataset.
@@
    Extract: We extract the linker Simplified Molecular-Input Line-Entry System (SMILES) strings for each high-performing MOF identified in Select step, eliminate those with more than three unique linker types, and assemble the remaining into a linker dataset.